# Implicit Bias of SignGD and Adam on Multiclass Separable Data

### v1

@@ 6 Implicit Bias of Adam
We generate snythetic multiclass separable data as follows: $k=5$ class centers are sampled from a standard normal distribution; within each class, data is sampled from normal distribution $\mathcal{N}(0,\sigma^{2}I),\sigma=0.1$. We set $d=25$, sample $5$ data points for each class, and ensure that margin is positive (thus data is separable). We run different algorithms to minimize CE loss using $\eta_{t}=\frac{\eta_{0}}{t^{a}}$ ($\eta_{0}=0.01$), where (based on our theorems) $a$ is set to $\nicefrac{{1}}{{2}}$, $\nicefrac{{1}}{{2}}$, and $\nicefrac{{2}}{{3}}$ for NGD, SignGD, and Adam, respectively. The stability constant $\epsilon$ for Adam is set from $\{0,10^{-6},10^{-7},10^{-8}\}$. We denote max-margin classifiers defined w.r.t. the 2-norm and the max-norm as ${\bm{V}}_{2}$ and ${\bm{V}}_{\infty}$ respectively. Fig. LABEL:fig:main_fig1 shows the following: (1) SignGD/Adam iterates favor the the max-norm margin over the 2-norm margin. The opposite is true for NGD (Figs. LABEL:fig:l2_margin,LABEL:fig:max_margin); (2) SignGD/Adam iterates correlate well with ${\bm{V}}_{\infty}$ whereas NGD correlates better with ${\bm{V}}_{2}$ (Figs. LABEL:fig:cor_inf,LABEL:fig:cor_v2); (3) Results are consistent for different (small) values of stability constant (curves nearly overlap).
@@ 7 Conclusion
We have characterized the implicit bias of SignGD, NSD, and Adam for multiclass separable data with CE loss, providing explicit rates for their margin maximization behavior. While these results establish fundamental theoretical guarantees, they are limited to linear models and separable data. Yet, they open several promising directions for future research: (1) Improving SignGD’s rates through appropriate momentum leveraging ideas from Ji et al. (2021); Wang et al. (2023) for NGD. (2) Extending the results to non-separable data (Ji & Telgarsky, 2019) and soft-label classification (Thrampoulidis, 2024) settings. (3) Extending the analysis to non-linear architectures like self-attention (Tarzanagh et al., 2023b, a; Vasudeva et al., 2024; Julistiono et al., 2024). Our treatment of CE loss through softmax properties in Sec. B is particularly relevant, as the implicit bias in self-attention is fundamentally driven by the softmax map. Moreover, self-attention naturally aligns with our multiclass setting since each token attends to multiple other tokens through softmax. (4) Investigating whether different implicit biases of SignGD/Adam to GD could theoretically explain empirical observations about their superior performance on heavy-tailed, imbalanced multiclass data (Kunstner et al., 2024). (5) From a statistical perspective, identifying specific scenarios where max-norm margin maximization leads to better generalization (building on initial investigations from (Salehi et al., 2019; Varma & Hassibi, 2024; Akhtiamov et al., 2024) on the impact of different regularization forms in linear classification, and recent findings from (Mohamadi et al., 2024) specifically related to infinity-norm regularization in neural networks).
@@ Acknowledgement
Appendix A Facts about CE loss and Softmax
Lemma A.1 is on the gradient of the cross-entropy loss. It will be used for showing the form of ${\mathcal{G}}({\bm{W}})$ in (7) lower bounds ${\left\|\nabla\mathcal{L}({\bm{W}})\right\|_{\max}}$ in Lemma B.1.
Lemma A.1 (Gradient).
@@ Appendix A Facts about CE loss and Softmax
Lemma A.2 is on the Taylor expansion of the loss. It will be used in showing the descent properties of SignGD (Lemma C.1) and Adam (Lemma D.5).
Lemma A.2 (Hessian).
@@
Lemma A.3 is used in bounding the second order term in the Taylor expansion of $\mathcal{L}({\bm{W}})$.
Lemma A.3.
For any $\bm{s}\in\Delta^{k-1}$ in the $k$-dimensional simplex, any index $c\in[k]$, and any ${\bm{v}}\in\mathbb{R}^{k}$ it holds:
${\bm{v}}^{\top}\left(\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\right){\bm{v}}\leq 4\,\|{\bm{v}}\|_{\infty}^{2}\,(1-s_{c})$
@@
By Cauchy-Schwartz,
${\bm{v}}^{\top}\left(\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\right){\bm{v}}$ $=\operatorname{vec}\big{(}\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\big{)}^{\top}\operatorname{vec}\big{(}{\bm{v}}{\bm{v}}^{\top}\big{)}$
$\leq\|\operatorname{vec}\big{(}\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\big{)}\|_{1}\|\operatorname{vec}\big{(}{\bm{v}}{\bm{v}}^{\top}\big{)}\|_{\infty}$
$\leq\|{\bm{v}}\|_{\infty}^{2}\,\|\operatorname{vec}\big{(}\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\big{)}\|_{1}\,.$
But,
$\|\operatorname{vec}\big{(}\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\big{)}\|_{1}$ $=\sum_{c\in[k]}s_{c}(1-s_{c})+\sum_{c^{\prime}\neq c\in[k]}s_{c}s_{c^{\prime}}$
$=2\sum_{c\in[k]}s_{c}(1-s_{c})$
where the last line uses $\sum_{c\neq c}s_{c^{\prime}}=1-s_{c}$.
The proof completes by applying Lemma A.4 to the above.
Lemma A.4 is used in the proof of Lemma A.3.
Lemma A.4.
@@
Appendix B Lemmas on $\mathcal{L}({\bm{W}})$ and ${\mathcal{G}}({\bm{W}})$
Lemma B.1 shows that ${\mathcal{G}}({\bm{W}})$ lower bounds the loss gradient and that the bound becomes tight as the loss approaches zero.
Lemma B.1 (${\mathcal{G}}({\bm{W}})$ as proxy to the loss-gradient norm).
Recall the margin $\gamma$ and the assumption $\|\bm{h}_{i}\|_{1}\leq B$. Then, for any ${\bm{W}}$ it holds that
$2B\cdot{\mathcal{G}}({\bm{W}})\geq{\left\|\nabla\mathcal{L}({\bm{W}})\right\|_{\rm{sum}}}\geq\gamma\cdot{\mathcal{G}}({\bm{W}})\,.$

### v4

@@ 6 Experiments
We generate snythetic multiclass separable data as follows: $k=10$ class centers are sampled from a standard normal distribution; within each class, data is sampled from normal distribution $\mathcal{N}(0,\sigma^{2}I),\sigma=0.1$. We set $d=25$, sample $50$ data points for each class, and ensure that margin is positive (thus data is separable). We run different algorithms to minimize CE loss using $\eta_{t}=\frac{\eta_{0}}{t^{a}}$ ($\eta_{0}=0.1$ for SignGD and NGD; $\eta_{0}=0.05$ for Spectral-GD and Muon), where (based on our theorems) $a$ is set to $\nicefrac{{1}}{{2}}$. We apply truncated SVD on the gradient and momentum for Spectral-GD and Muon respectively. Data margins w.r.t. different norms are found via CVXPY [diamond2016cvxpy]. We denote max-margin classifiers defined w.r.t. the 2-norm, the max-norm, and the spectral-norm as ${\bm{V}}_{2}$, $\bm{V}$$_{\infty}$, and ${\bm{V}}_{\operatorname{spec}}$, respectively. Based on the margin-gap results in Figure 1, we observe that SignGD, NGD, and Spectral-GD favor max-norm, 2-norm, and spectral-norm margin respectively. Besides this, the behavior of Muon is very similar to that of Spectral-GD (in agreement with our theories). Figure 2 further confirms that the iterates of these algorithms correlate well with the corresponding max margin separators. Experiments on Signum, NMD-GD, and Adam are provided in App. A and App. G.
Two-layer Neural Network
We extend the experiments to the non-linear classification setting using the cross-entropy loss. We sample 100 data points from each of the 10 classes of the MNIST dataset [lecun2002gradient]. The model is a two-layer neural network with the hidden dimension being 100 (the first and second layer weights are denoted as $\bm{V}$} and $\bm{W}$} respectively). The training is done in two ways: (a) Train the first-layer weight with the second-layer weight fixed and (b) Train both the first and second-layer weights. For options (a) and (b), the respective spectral-norm margins of the overall network are defined as
[…]
where $\sigma(\cdot)$ is the sigmoid function. In Figure 3, we track the quantities $\gamma_{a}^{{\bm{V}}_{t}}$ (Figure 3(a) and 3(b)) and $\gamma_{a}^{{\bm{V}}_{t},{\bm{W}}_{t}}$ (Figure 3(c) and 3(d)) as a function of the iteration counter $t$. For both definitions, we observe that the spectral-norm margin of the iterates of Spectral-GD and Muon grow faster than other algorithms. Hence, the observations in the linear settings can also hold in the non-linear settings.
(a) SignGD
(b) NGD
(c) Spectral-GD
(d) Muon
Figure 1: (a) We normalize the iterates of SignGD w.r.t. the max-norm (denoted as $\bar{{\bm{W}}}_{t}$), compute the margin (denoted as $\gamma_{\bar{{\bm{W}}}_{t}}$), then plot its difference to data margins $\gamma_{\lVert\cdot\rVert_{\infty}}$, $\gamma_{\lVert\cdot\rVert_{2}}$, and $\gamma_{\left|\left|\left|\cdot\right|\right|\right|_{\infty}}$ (note that the margin difference is further divided by the corresponding data margin for comparisons). SignGD favors the margin defined w.r.t. the max-norm. (b, c, and d) Same as (a) with SignGD (max-norm) replaced by NGD (2-norm), Spectral-GD (spectral-norm), and Muon (spectral-norm), respectively. NGD favors the 2-norm margin, while Spectral-GD and Muon favor the spectral-norm margin.
(a) SignGD
(b) NGD
(c) Spectral-GD
(d) Muon
Figure 2: (a) Correlations between the iterates of SignGD (${\bm{W}}_{t}$) and max margin separators $\bm{V}$$_{\infty}$, ${\bm{V}}_{2}$, and ${\bm{V}}_{\operatorname{spec}}$ against iterations (correlation defined as $\frac{\langle{\bm{W}},{\bm{V}}\rangle}{\lVert{\bm{W}}\rVert_{2}\lVert{\bm{V}}\rVert_{2}}$). (b, c, and d) Same as (a) with SignGD replaced by NGD, Spectral-GD, and Muon, respectively. SignGD and NGD correlate well with $\bm{V}$$_{\infty}$ and ${\bm{V}}_{2}$, respectively, while Spectral-GD and Muon correlate well with ${\bm{V}}_{\operatorname{spec}}$.
(a) Single Layer($\gamma_{a}^{{\bm{V}}}$)
(b) Single Layer($\gamma_{a}^{{\bm{V}}}$)
(c) Joint Training($\gamma_{b}^{{\bm{V}},{\bm{W}}}$)
(d) Joint Training($\gamma_{b}^{{\bm{V}},{\bm{W}}}$)
Figure 3: (a) Spectral-norm margin $\gamma_{a}^{{\bm{V}}_{t}}$ as a function of $t$ for SignGD, NGD, and Spectral-GD. (b) Same as (a) with algorithms replaced by Signum, NMD-GD, and Muon, respectively. (c) Spectral-norm margin $\gamma_{b}^{{\bm{V}}_{t},{\bm{W}}_{t}}$ as a function of $t$ for SignGD, NGD, and Spectral-GD. (d) Same as (c) with algorithms replaced by Signum, NMD-GD, and Muon, respectively.
7 Related Works
Starting with GD, the foundational result by soudry2018implicit showed that gradient descent optimization of logistic loss on linearly separable data converges in direction to the $L_{2}$ max-margin classifier at a rate $O(1/\log(t))$. Contemporaneous work by ji2019implicit generalized this by relaxing the data separability requirement. ji2020gradient later connected these findings to earlier work on regularization paths of logistic loss minimization [Rosset2003MarginML], which enabled extensions to other loss functions (e.g., those with polynomial tail decay). More recently, wu2024implicit extends these results to the large step size regime with the same $O(1/\log(t))$ rate. The relatively slow convergence rate to the max-margin classifier motivated investigation into adaptive step-sizes. nacson2019convergence showed that NGD with decaying step-size $\eta_{t}=1/\sqrt{t}$ achieves $L_{2}$-margin convergence at rate $O(1/\sqrt{t})$. This rate was improved to $O(1/t)$ by ji2021characterizing using constant step-sizes, and further to $O(1/t^{2})$ through a specific momentum formulation [ji2021fast]. Besides linear classifications, implicit bias of GD has been studied for least squares [gunasekar2017implicit, gunasekar2018characterizing, azizan2021stochastic], homogeneous [lyu2019gradient, ji2020directional, wu2024large] and non-homogeneous neural networks [cai2025implicit], as well as matrix factorization [gunasekar2017implicit]; see vardi2023implicit for a survey.
All the above mentioned works focus almost exclusively on binary classification. The noticeable gap in analysis of multiclass classification in most existing literature is highlighted by seli, and more recently emphasized by ravi2024implicit, who extended the implicit bias result of soudry2018implicit to multiclass classification for losses with exponential tails, including CE, multiclass exponential, and PairLogLoss. Their approach leverages a framework of wang2024unified that allows multiclass losses and separability conditions to be written in margin-based forms similar to binary cases. However, these works only focus on GD with the $L_{2}$-geometry. In this work, we consider a wide range of algorithms with different geometries for multiclass classification.
Beyond GD, gunasekar2018characterizing and nacson2019convergence showed that steepest descent w.r.t. entry-wise p-norms yields updates that in the limit maximize the margin w.r.t the same norm. sun2022mirror, sun2023unified showed that the iterates of mirror descent with the potential function chosen as the p-th power of the p-norm converge to the classifier that maximizes the margin w.r.t. the p-norm. In both cases, the convergence rate is slow at $O(1/\log(t))$. wang2023faster further improved the rates for both steepest descent and mirror descent when $p\in(1,2]$. Note that all these results apply only to the exponential loss. More recently, tsilivis2024flavors showed that the iterates of steepest descent algorithms converge to a KKT point of a generalized margin maximization problem in homogeneous neural networks. Moreover, the implicit bias of Adam (with or without the stability constant) has been studied in both linear and non-linear settings. wang2021implicit demonstrated the normalized iterates of Adam (with non-negligible stability constant) converge to a KKT point of a $L_{2}$-margin maximization problem for homogeneous neural networks. zhang2024implicit studied the implicit bias of Adam without the stability constant on (linearly) binary separable data. They showed that unlike GD, the Adam’s iterates converge to a solution that maximizes the margin w.r.t the $L_{\infty}$-norm. The study of excluding the stability constant is also the focus of another recent work on the implicit bias of AdamW [xie2024implicit], where the authors again establish that convergence aligns with the $L_{\infty}$ geometry.
@@ 8 Conclusion
We have characterized the margin convergence rates of Spectral-GD and Muon for multiclass linear separable data. Given they are special cases of NSD and NMD w.r.t the spectral norm, the analysis is done on a wider scale by studying NSD/NMD w.r.t any entry-wise or Schatten p-norms. Thus, the rates also hold for optimizers of other geometries, such as the sign-descent (max-norm) or gradient-descent (2-norm) family. We further extend the analysis to Adam using the same framework. Future directions include removing the factor-$d$ from the bound of NMD, obtaining a tighter convergence rate for Adam, and studying other related algorithms such as Shampoo that involves non-diagonal preconditioners. It is also important to extend our results to (multiclass) non-separable settings [ntp] and nonlinear models such as diagonal neural nets [pesme2021implicit], self-attention mechanisms [tarzanagh2023maxmargin, ataee2023max, tarzanagh2023transformers, deora2024implicit, julistiono2024optimizing] and homogeneous neural nets [lyu2019gradient, tsilivis2024flavors, cai2025implicit], helping further bridge the gap to deep learning practices. Finally, from a complementary statistical perspective, future work could seek identifying specific scenarios where margin maximization with respect to norms other than Frobenius leads to better generalization (extending a long line of prior works, e.g., [salehi2019impact, deng2019model, koehler2021uniform, muthukumar2020classification, wang2021benign, donhauser2022fast, taheri2023generalization, varma2024benefits, akhtiamov2024regularized, tsigler2025benign]).
@@ Acknowledgement
Appendix A Additional Experiments
We present additional experiments on Signum and NMD-GD in this section. Based on Figure 4, their margin convergence properties are very similar to those of SignGD and NGD respectively (see Sec. 6 in the main text for the experimental setup).
(a) Signum
(b) Signum
(c) NMD-GD
(d) NMD-GD
Figure 4: Implicit bias of Signum and NMD-GD on multiclass separable data. (a) Relative margin gap of Signum’s iterates against iterations. (b) Correlation of Signum’s iterates to $\bm{V}$$_{\infty}$, ${\bm{V}}_{2}$, and ${\bm{V}}_{\operatorname{spec}}$ against iterations. See Figure 1 and 2 for the definitions of relative margin and correlation. (c) and (d) Same as (a) and (b) with Signum replaced by NMD-GD.
Appendix B Facts about CE loss and Softmax
Lemma 9 is on the gradient of the cross-entropy loss. It will be used for showing the form of ${\mathcal{G}}({\bm{W}})$ in (8) lower bounds $\lVert\nabla\mathcal{L}({\bm{W}})\rVert$ in Lemma 16.
Lemma 9 (Gradient).
@@ Appendix B Facts about CE loss and Softmax
Lemma 10 is on the Taylor expansion of the loss. It will be used in showing the descent properties of NSD and NMD.
Lemma 10 (Hessian).
@@
We prove the relationships in (7), which are useful for unifying the analysis of entry-wise and Schatten norms.
Lemma 11.
For any matrix $\bm{A}\in\mathbb{R}^{m\times n}$ and any entry-wise or Schatten p-norm $\lVert\cdot\rVert$ with $p\geq 1$, it holds that
${\left\|\bm{A}\right\|_{\max}}\leq\lVert\bm{A}\rVert\leq{\left\|\bm{A}\right\|_{\rm{sum}}}\,.$
@@
[…]
It is also well-known that
[…]
where the inequality follows by selecting $\bm{u}=\texttt{sign}(\bm{A}[i^{\prime},j^{\prime}])\cdot\bm{e}_{i^{\prime}}$ and ${\bm{v}}=\bm{e}_{j^{\prime}}$ for $(i^{\prime},j^{\prime})$ such that $|\bm{A}[i^{\prime},j^{\prime}]|={\left\|\bm{A}\right\|_{\max}}$ and $\bm{e}_{i^{\prime}},\bm{e}_{j^{\prime}}$ corresponding basis vectors.
Using this together with duality, it also holds that
[…]
This follows from the following sequnece of inequalities
[…]
where the first inequality follows from generalized Cauchy-Scwhartz and the second inequality by (16). ∎
Lemma 12 is used in bounding the second order term in the Taylor expansion of $\mathcal{L}({\bm{W}})$.
Lemma 12.
For any entry-wise or Schatten p-norm $\lVert\cdot\rVert$ with $p\geq 1$, any $\bm{s}\in\Delta^{k-1}$ in the $k$-dimensional simplex, any index $c\in[k]$, and ${\bm{v}}\in\mathbb{R}^{k}$, it holds that
${\bm{v}}^{\top}\left(\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\right){\bm{v}}\leq 4(1-s_{c})\lVert{\bm{v}}{\bm{v}}^{T}\rVert.$
Proof.
See main text. ∎
Lemma 13 is used in the proof of Lemma 12.
Lemma 13.
@@
Lemma 14.
For any $\bm{s}\in\Delta^{k-1}$ in the $k$-dimensional simplex, any index $c\in[k]$, any $\bm{\Delta}\in\mathbb{R}^{k\times d}$, and any $\bm{h}\in\mathbb{R}^{d}$, it holds:
$\bm{h}^{\top}\bm{\Delta}^{\top}\left(\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\right)\bm{\Delta}\bm{h}\leq 4B^{2}\lVert\bm{\Delta}\rVert^{2}\,(1-s_{c})\,.$